# A Practical Framework for Estimating the Repetition Likelihood of Fast Radio Bursts from Spectral Morphology

### v1

@@ Abstract
The repeating behavior of fast radio bursts (FRBs) is regarded as a key clue to understanding their physical origin, yet reliably distinguishing repeaters from apparent non-repeaters with current observations remains challenging. Here we propose a physically interpretable and practically quantifiable classification framework based on spectral morphology. Using dimensionality reduction, clustering, and feature-importance analysis, we identify the spectral running $r$ and spectral index $\gamma$ as the most critical parameters for distinguishing repeaters from apparent non-repeaters in the CHIME/FRB sample. In the $\gamma$-$r$ space, repeaters preferentially occupy regions with steeper, narrower-band spectra, whereas non-repeaters cluster in flatter, broader-band regions, resulting in a clear density separation. We further construct an empirical probability map in the $\gamma$-$r$ space, showing a clear gradient of repetition likelihood, from $\sim$65% in the high-repetition region to $\sim$5% in the low-repetition region. Combining this with Gaussian Mixture Model posterior analysis, we identify several apparent non-repeaters with high inferred repetition probability, recommending them as priority targets for future monitoring. This framework provides a simple and generalizable tool for assessing repeatability in the CHIME/FRB sample and highlights the diagnostic power of spectral morphology in unveiling FRB origins.
@@ 1 Introduction
In recent years, numerous studies have attempted to distinguish repeaters from apparent non-repeaters using supervised machine learning approaches (Luo et al., 2023; Sharma & Rajpaul, 2024; Kharel et al., 2025) and unsupervised (Chen et al., 2022; Yang et al., 2023; Zhu-Ge et al., 2023; García et al., 2024; Sun et al., 2025; Qiang et al., 2025; Júnior et al., 2025; Liu et al., 2025). Generally, the supervised approaches face the primary challenge of avoiding erroneous training and selecting representative, well-labeled datasets for robust model development. Given current observational limitations, inappropriate training sets or biases can lead to misleading classifications, necessitating caution in their application.
@@
To address these challenges, we propose a physically interpretable and statistically robust framework that estimates the probability of an FRB exhibiting repeater-like properties, thereby providing a probabilistic metric to inform follow-up strategies. In this work, we perform a systematic analysis of multidimensional observational parameters using an extended CHIME/FRB sample (combining the first CHIME/FRB catalog (CHIME/FRB Collaboration et al., 2021) with newly reported repeaters from CHIME/FRB Collaboration et al. (2023)). The structure of our paper is as follows: In Section 2, we introduce the extended CHIME/FRB sample and the methods used in this work. Section 3 presents the clustering results and the discussions. In Section 4, we summarize our work.
2 DATA AND ALGORITHM PARAMETERS
2.1 Data Sample Selection
@@ 2.2 FRB Feature Parameters
3 RESULTS AND DISCUSSION
3.1 t-SNE Results
@@ 3 RESULTS AND DISCUSSION
While the Repeater-like cluster is dominated by repeating FRBs, it also includes some apparent non-repeaters. Conversely, the Nonrepeater-like cluster consists primarily of apparent non-repeaters but contains a few repeaters as well. Interestingly, some repeating bursts assigned to both clusters originate from the same repeating source. Those repeating bursts in the Nonrepeater-like cluster are rare and tend to exhibit broader bandwidths, narrower pulse widths, and higher energies than typical Repeater-like bursts, representing atypical repeating activity. Such cases demonstrate that repeaters can span a broad parameter space and occasionally resemble typical non-repeaters, reinforcing the view that the two clusters reflect dominant but overlapping similarity patterns in the nine-dimensional parameter space, possibly shaped by different physical mechanisms or environments. Moreover, the distribution of repeaters within the Repeater-like cluster itself appears inhomogeneous, suggesting the presence of substructures or subclusters that merit detailed investigation with larger samples in future studies.
@@ 3.2 Feature Importance
3.3 $\gamma$-$r$ Parameter Distributions
3.3.1 $\gamma$-$r$ Space Number Density Distribution
@@
Overall, the $\gamma$-$r$ space exhibits both continuity and multimodality. The significant separation between the density peaks of repeaters and non-repeaters suggests that repetition correlates with specific spectral properties, potentially reflecting intrinsic differences in emission or propagation mechanisms. However, current evidence remains insufficient to conclusively determine whether repeaters and non-repeaters originate from fundamentally distinct source classes. A more cautious interpretation is that the observed separation represents two characteristic regimes, possibly reflecting extremes of a continuous spectrum of behaviors produced by a single population or physical process operating under different conditions, as discussed in Section 3.3.3.
3.3.2 $\gamma$-$r$ Space Correlation with Bandwidth, Pulse Width, and Burst Rate
@@
3.3.3 Physical Interpretation within Geometric Frameworks
@@
Based on the above analysis, a more plausible interpretation is that the separation of density centers and the correlation between $\gamma$-$r$ distributions indicate that repeating and non-repeating FRBs may not arise from two fundamentally distinct progenitor mechanisms. Instead, they more likely represent two manifestations of the same underlying mechanism or source class under different geometric and physical conditions, i.e., two extreme cases within a continuous distribution. The rotating polar cap scenario (Beniamini & Kumar, 2025; Luo et al., 2025) provides a coherent framework for interpreting these results. In this framework, whether an FRB appears as repeating or non-repeating, along with its associated spectral and temporal characteristics, is mainly governed by the viewing geometry of the line of sight relative to the magnetar’s emission cone. In the aligned-rotator geometry ($\alpha\ll\rho$), where $\alpha$ is the magnetic inclination angle and $\rho$ is the half-opening angle of the emitting beam in the observer’s frame, the line of sight remains within a fixed emission region, so the physical conditions vary little, leading to pulse widths set mainly by the intrinsic burst timescale and spectra centered around a stable frequency. Such emission conditions can be naturally sustained in magnetars residing in binary systems (Zhang & Hu, 2025). This produces bursts with narrower spectra, longer widths, and higher repetition rates, corresponding to the clustering of repeating FRBs at $r\ll 0$ and relatively large $\gamma$ in Figure 3. In contrast, in the misaligned-rotator geometry ($\alpha\gg\rho$), the line of sight samples a wide range of surface regions as the beam sweeps by, causing large variations in emission conditions and yielding bursts with sweep-limited widths, broadened spectra, and lower apparent repetition rates. These correspond to the clustering of non-repeating FRBs around $r\sim 0$ and relatively smaller $\gamma$ in the $\gamma$-$r$ space. This geometric framework naturally explains the pronounced separation yet partial overlap of the density centers of repeating and non-repeating FRBs observed in the $\gamma$-$r$ plane.
@@
The geometry framework offers a potential explanation for the observed differences between repeating and apparently non-repeating FRBs, although doing so hinges on restrictive geometric parameters. In particular, to simultaneously reproduce the nearly flat polarization position angles (PAs) commonly observed in both repeating and apparently non-repeating FRBs (Michilli et al., 2018; CHIME/FRB Collaboration et al., 2019; Luo et al., 2020; Hilmarsson et al., 2021b, a; Xu et al., 2022a; Jiang et al., 2022; Sand et al., 2022; Nimmo et al., 2022; Zhang et al., 2023b; Mckinven et al., 2023; Kumar et al., 2023; Pandhi et al., 2024; Faber et al., 2024; Liu et al., 2025; Xie et al., 2025; Ng et al., 2025), the rotating vector model (Radhakrishnan & Cooke, 1969) requires that at least one of the following conditions be satisfied: an exceptionally long rotational period $P\gg\Delta t_{\mathrm{WS}}$, nearly aligned rotational and magnetic axes $\sin(\alpha)\sim 0$, or a very small impact angle $|\beta|\sim 0$. The period requirement is expected for repeating FRBs, as current statistical analyses of large samples from some hyperactive repeating FRBs have excluded the existence of periods in the millisecond-to-second range (Li et al., 2021; Xu et al., 2022a; Niu et al., 2022; Zhang et al., 2023b; Wang et al., 2025; Zhou et al., 2025b). In contrast, such a long period is not anticipated for non-repeating FRBs, since their pulse width $\Delta t_{\mathrm{WS}}$ relies on the rotational cut-off. The second condition of aligned spin and magnetic axes is also plausible for repeating FRBs. However, non-repeating FRBs further require $\alpha\gg\rho$. Under this premise, both $\alpha$ and $\rho$ must be very small, which constitutes a highly restrictive constraint. The third requirement on the impact angle applies equally to both types of FRBs. An alternative explanation for the observed disparities between the repeating and non-repeating FRBs may lie in intrinsic variations within the emission mechanism itself, which also governs the resulting spectral morphology.
@@
We note that in Figure 3, a small subset of repeaters falls within the primary density peak of apparent non-repeaters. These bursts exhibit broader spectra and shorter durations, and generally originate from only a few sporadic events of a given repeater. Their properties differ markedly from those of the majority of typical repeating bursts. This suggests that the emission region or emission mechanism of repeaters may not always remain stable and consistent. Instead, different physical processes, such as changes in the emission site or beaming angle, may occasionally produce short-duration, broadband, and potentially high-energy bursts that resemble those of apparently non-repeating sources (Kirsten et al., 2024). In addition, such “outlier” events may also be explained by low-probability processes external to the emission source. For example, propagation effects such as plasma lensing can, on rare occasions, significantly modify the observed spectra of bursts (Cordes et al., 2017; Sobacchi et al., 2021; Kumar et al., 2024). However, their occurrence rate is expected to be extremely low, and thus they can only account for a small number of anomalous events. On the other hand, statistical fluctuations in a limited sample can naturally produce a few events located at the tails of the underlying parameter distribution, making these outliers statistically plausible. Consequently, the continuous distribution and density structure differences observed in the $\gamma$-$r$ space may not only arise from geometric effects within a single theoretical framework, but also reflect continuous manifestations of the same class of sources in different physical states. This provides important observational clues for probing the origins and emission mechanisms of FRBs.
3.4 $\gamma$-$r$ Framework for Repetition Probability Estimation
@@
In this section, we construct an empirical probability map in the $\gamma$-$r$ space to estimate the likelihood of FRBs exhibiting repeating behavior. In Figure 4, we draw dashed lines at the median values of $r$ and $\gamma$ for the entire FRBs, dividing the parameter space into four quadrants. This division approximately corresponds to the natural boundary between the density peaks of non-repeaters and repeaters shown in Figure 3, reflecting that certain spectral morphologies are more likely to produce repeating FRBs. Each region represents a typical combination of spectral properties, and we further calculate the fraction of repeaters within each region as an approximate estimate of the repetition probability: 0.65 (I), 0.36 (II), 0.28 (III), and 0.05 (IV). This pronounced gradient demonstrates a clear dependence of FRB repeatability on the $\gamma$-$r$ parameter space. In particular, the stark contrast between the high-repetition Region I and the nearly non-repeating Region IV further quantifies the association between spectral morphology and repeat behavior. This result is consistent with the previous analysis based on density contours, reinforcing spectral morphology as a potential indicator for distinguishing repeaters.
@@
Figure 4: Distribution of repeaters (red) and apparent non-repeaters (blue) in the $\gamma$-$r$ parameter space. The background shading indicates regions of different inferred repetition probabilities, with annotated values representing the fraction of known repeaters in each quadrant. FRB 20201124A and FRB 20121102A are highlighted as cyan and yellow diamonds, respectively.
Additionally, in Figure 4, we highlight two well-studied and highly active repeating FRB sources: FRB 20121102A and FRB 20201124A. Both have been monitored extensively with high sensitivity by the Arecibo telescope (Jahns et al., 2023) and FAST (Li et al., 2021; Xu et al., 2022b), yielding records of thousands of bursts, thereby providing critical insights into the origins of FRBs. Notably, both sources fall within the high-repetition-probability Region I defined in our framework, serving as compelling evidence for the robustness and predictive usefulness of this spectral partitioning. Except for a small number of atypical repeating bursts, this framework successfully assigns the vast majority of repeaters to the high-probability Region I. Overall, this probability map directly links FRB repeatability to spectral morphology parameters, offering a simple, physically interpretable prediction framework independent of prior labeling. For future FRBs newly detected by CHIME, their positions in this map can be readily determined using only $r$ and $\gamma$, enabling a preliminary estimation of their likelihood to exhibit repeating FRBs. This provides a practical and accessible tool for rapidly assessing repetition probability and prioritizing monitoring targets in future observations. Compared with machine learning classification methods that rely on multi-parameter dimensionality reduction and clustering, this physically based partitioning is faster, more intuitive, and yields robust, generalizable results.
@@
We further combine the t-SNE results with the $\gamma$-$r$ parameter map to identify a set of FRB sources that have not yet been observed to repeat but exhibit statistical characteristics closely resembling those of known repeaters. Specifically, we first selected all non-repeating FRBs classified within the Repeater-like cluster and categorized them according to their quadrant positions in the $\gamma$-$r$ parameter space, with particular attention to those located in the high-repetition-probability Region I. In addition, we incorporated a Bayesian posterior probability assessment based on GMM fitting to model the probability density of repeaters and non-repeaters in the $\gamma$-$r$ parameter space, thereby estimating the posterior probability that each FRB belongs to the repeater population. In this framework, the posterior probability for a given FRB point $\mathbf{x}=(\gamma,r)$ to be classified as a repeater is computed using Bayes’ theorem,
@@
4 Conclusion
In this Letter, we analyzed the repeatability characteristics of FRBs using an extended CHIME/FRB sample, integrating dimensionality reduction, clustering, feature-importance evaluation, and spectral-parameter-based partitioning. We further propose a simple and generalizable method that relies solely on spectral parameters to classify FRBs by their repeatability preference and to effectively estimate their repetition probabilities. The main conclusions are as follows:

### v2

@@ Abstract
The repeating behavior of fast radio bursts (FRBs) is regarded as a key clue to understanding their physical origin, yet reliably distinguishing repeaters from apparent non-repeaters with current observations remains challenging. Here we propose a physically interpretable and practically quantifiable classification framework based on spectral morphology. Using dimensionality reduction, clustering, and feature-importance analysis, we identify the spectral running $r$ and spectral index $\gamma$ as the most critical parameters for distinguishing repeaters from apparent non-repeaters in the CHIME/FRB sample. In the $\gamma$-$r$ space, repeaters preferentially occupy regions with steeper, narrower-band spectra, whereas non-repeaters cluster in flatter, broader-band regions, resulting in a clear density separation. We further construct a probability map in the $\gamma$-$r$ space based on Gaussian Mixture Model posterior analysis, revealing a clear gradient of repetition likelihood from $\sim$67% in the high-repetition region to $\sim$7% in the low-repetition region. This model also identifies several apparent non-repeaters with high inferred repetition probability, highlighting them as priority targets for future monitoring. This framework provides a simple and generalizable tool for assessing repeatability in the CHIME/FRB sample and highlights the diagnostic power of spectral morphology in unveiling FRB origins.
@@ I Introduction
In recent years, numerous studies have attempted to distinguish repeaters from apparent non-repeaters using supervised machine learning approaches (Luo et al., 2023; Sharma & Rajpaul, 2024; Kharel et al., 2025) and unsupervised (Chen et al., 2022; Yang et al., 2023; Zhu-Ge et al., 2023; García et al., 2024; Sun et al., 2025; Qiang et al., 2025; Liu et al., 2025a; Madheshwaran et al., 2025; Júnior et al., 2026). Generally, the supervised approaches face the primary challenge of avoiding erroneous training and selecting representative, well-labeled datasets for robust model development. Given current observational limitations, inappropriate training sets or biases can lead to misleading classifications, necessitating caution in their application.
@@
To address these challenges, we propose a physically interpretable and statistically robust framework that estimates the probability of an FRB exhibiting repeater-like properties, thereby providing a probabilistic metric to inform follow-up strategies. In this work, we perform a systematic analysis of multidimensional observational parameters using an extended CHIME/FRB sample (combining the first CHIME/FRB catalog (CHIME/FRB Collaboration et al., 2021) with newly reported repeaters from CHIME/FRB Collaboration et al. (2023)). The structure of our paper is as follows: In Section II, we introduce the extended CHIME/FRB sample and the methods used in this work. Section III presents the clustering results and the discussions. In Section IV, we summarize our work.
II DATA AND ALGORITHM PARAMETERS
II.1 Data Sample Selection
@@ II.2 FRB Feature Parameters
III RESULTS AND DISCUSSION
III.1 t-SNE Results
@@ III RESULTS AND DISCUSSION
While the Repeater-like cluster is dominated by repeating FRBs, it also includes some apparent non-repeaters. Conversely, the Nonrepeater-like cluster consists primarily of apparent non-repeaters but contains a few repeaters as well. Interestingly, some repeating bursts assigned to both clusters originate from the same repeating source. Those repeating bursts in the Nonrepeater-like cluster are rare and tend to exhibit broader bandwidths, narrower pulse widths, and higher energies than typical Repeater-like bursts, representing atypical repeating activity. Such cases demonstrate that repeaters can span a broad parameter space and occasionally resemble typical non-repeaters, reinforcing the view that the two clusters reflect distinct but overlapping similarity patterns in the nine-dimensional parameter space, possibly shaped by different physical mechanisms or environments. Moreover, the distribution of repeaters within the Repeater-like cluster itself appears inhomogeneous, suggesting the presence of substructures or subclusters that merit detailed investigation with larger samples in future studies.
@@ III.2 Feature Importance
III.3 $\gamma$-$r$ Parameter Distributions
III.3.1 $\gamma$-$r$ Space Number Density Distribution
@@
Overall, the $\gamma$-$r$ space exhibits both continuity and multimodality. The significant separation between the density peaks of repeaters and non-repeaters suggests that repetition correlates with specific spectral properties, potentially reflecting intrinsic differences in emission or propagation mechanisms. However, current evidence remains insufficient to conclusively determine whether repeaters and non-repeaters originate from fundamentally distinct source classes. A more cautious interpretation is that the observed separation represents two characteristic regimes, possibly reflecting extremes of a continuous spectrum of behaviors produced by a single population or physical process operating under different conditions, as discussed in Section III.3.3.
III.3.2 $\gamma$-$r$ Space Correlation with Bandwidth, Pulse Width, and Burst Rate
@@
III.3.3 Physical Interpretation within Geometric Frameworks
@@
Based on the above analysis, FRBs in the $\gamma$-$r$ space exhibit a positive correlation between repetition rate and pulse width and a negative correlation with bandwidth. These trends define a continuous, monotonic transition from non-repeaters to repeaters, suggesting that the two populations may not arise from fundamentally distinct progenitor mechanisms. Instead, they more likely represent two manifestations of the same underlying mechanism or source class under different geometric and physical conditions, i.e., two extreme cases within a continuous distribution. The rotating polar cap scenario (Beniamini & Kumar, 2025a; Luo et al., 2025) provides a coherent framework for interpreting these results. In this framework, whether an FRB appears as repeating or non-repeating, along with its associated spectral and temporal characteristics, is mainly governed by the viewing geometry of the line of sight relative to the magnetar’s emission cone. In the aligned-rotator geometry ($\alpha\ll\rho$), where $\alpha$ is the magnetic inclination angle and $\rho$ is the half-opening angle of the emitting beam in the observer’s frame, the line of sight remains within a fixed emission region, so the physical conditions vary little, leading to pulse widths set mainly by the intrinsic burst timescale and spectra centered around a stable frequency. Such emission conditions can be naturally sustained in magnetars residing in binary systems (Zhang & Hu, 2025). This produces bursts with narrower spectra, longer widths, and higher repetition rates, corresponding to the clustering of repeating FRBs at $r\ll 0$ and relatively large $\gamma$ in Figure 3. In contrast, in the misaligned-rotator geometry ($\alpha\gg\rho$), the line of sight samples a wide range of surface regions as the beam sweeps by, causing large variations in emission conditions and yielding bursts with sweep-limited widths, broadened spectra, and lower apparent repetition rates. These correspond to the clustering of non-repeating FRBs around $r\sim 0$ and relatively smaller $\gamma$ in the $\gamma$-$r$ space. This geometric framework naturally explains the pronounced separation yet partial overlap of the density centers of repeating and non-repeating FRBs observed in the $\gamma$-$r$ plane (see Figure 3).
@@
The geometry framework offers a potential explanation for the observed differences between repeating and apparently non-repeating FRBs, although doing so hinges on restrictive geometric parameters. In particular, to simultaneously reproduce the nearly flat polarization position angles (PAs) commonly observed in both repeating and apparently non-repeating FRBs (Michilli et al., 2018; CHIME/FRB Collaboration et al., 2019; Luo et al., 2020; Hilmarsson et al., 2021b, a; Xu et al., 2022a; Jiang et al., 2022; Sand et al., 2022; Nimmo et al., 2022; Zhang et al., 2023b; Mckinven et al., 2023; Kumar et al., 2023; Pandhi et al., 2024; Faber et al., 2024; Liu et al., 2025b; Xie et al., 2025; Ng et al., 2025), the rotating vector model (Radhakrishnan & Cooke, 1969) requires that at least one of the following conditions be satisfied:
(1) an exceptionally long rotational period $P\gg\Delta t_{\mathrm{WS}}$;
(2) nearly aligned rotational and magnetic axes $\sin(\alpha)\sim 0$;
(3) a very small impact angle $|\beta|\sim 0$.
For condition (1), the requirement of a long rotational period is expected for repeating FRBs, as current statistical analyses of large samples from some hyperactive repeating FRBs have excluded the existence of periods in the millisecond-to-second range (Li et al., 2021; Xu et al., 2022a; Niu et al., 2022; Zhang et al., 2023b; Wang et al., 2025; Zhou et al., 2025b; Zhang et al., 2025). In contrast, such a long period is not anticipated for non-repeating FRBs, since their pulse width $\Delta t_{\mathrm{WS}}$ relies on the rotational cut-off. For condition (2), near alignment between the spin and magnetic axes is also plausible for repeating FRBs. However, for non-repeating FRBs this condition further requires $\alpha\gg\rho$. Under this premise, both $\alpha$ and $\rho$ must be very small, which constitutes a highly restrictive constraint. For condition (3), the requirement of a very small impact angle applies equally to both types of FRBs. Beyond the geometric conditions discussed above, an alternative explanation for the observed disparities between the repeating and non-repeating FRBs may lie in intrinsic variations within the emission mechanism itself, which also governs the resulting spectral morphology.
@@
We note that in Figure 3, a small subset of repeaters falls within the primary density peak of apparent non-repeaters. These bursts exhibit broader spectra and shorter durations, and generally originate from only a few sporadic events of a given repeater. Their properties differ markedly from those of the majority of typical repeating bursts. This suggests that the emission region or emission mechanism of repeaters may not always remain stable and consistent. Instead, different physical processes, such as changes in the emission site or beaming angle, may occasionally produce short-duration, broadband, and potentially high-energy bursts that resemble those of apparently non-repeating sources (Kirsten et al., 2024). In addition, such “outlier” events may also be explained by low-probability processes external to the emission source. For example, propagation effects such as plasma lensing can, on rare occasions, significantly modify the observed spectra of bursts (Cordes et al., 2017; Sobacchi et al., 2021; Kumar et al., 2024). However, their occurrence rate is expected to be extremely low, and thus they can only account for a small number of anomalous events. On the other hand, statistical fluctuations in a limited sample can naturally produce a few events located at the tails of the underlying parameter distribution, making these outliers statistically plausible. Consequently, the continuous distribution and density structure differences observed in the $\gamma$-$r$ space may not only arise from geometric effects within a single theoretical framework, but also reflect continuous manifestations of the same class of sources in different physical states (Beniamini & Kumar, 2025b). This provides important observational clues for probing the origins and emission mechanisms of FRBs.
III.4 $\gamma$-$r$ Framework for Repetition Probability Estimation
@@
In this section, we construct an empirical probability map in the $\gamma$-$r$ space to estimate the likelihood of FRBs exhibiting repeating behavior. In Figure 4, we draw dashed lines at the median values of $r$ and $\gamma$ for the entire FRBs, dividing the parameter space into four quadrants. This division approximately corresponds to the natural boundary between the density peaks of non-repeaters and repeaters shown in Figure 3, reflecting that certain spectral morphologies are more likely to produce repeating FRBs.
@@
Figure 4: Distribution of repeaters (red) and apparent non-repeaters (blue) in the $\gamma$-$r$ parameter space. The background shading indicates regions of different inferred repetition probabilities, with annotated values representing the mean posterior probabilities derived from the Bayesian GMM analysis. FRB 20201124A and FRB 20121102A are highlighted as cyan and yellow diamonds, respectively.
@@
To further quantify this distinction, we incorporated a Bayesian posterior probability assessment based on GMM fitting to model the probability density of repeaters and non-repeaters in the $\gamma$-$r$ parameter space, thereby estimating the posterior probability that each FRB belongs to the repeater population. In this framework, the posterior probability for a given FRB point $\mathbf{x}=(\gamma,r)$ to be classified as a repeater is computed using Bayes’ theorem,
@@
Based on these posterior probabilities, we computed the average repetition probability within each region defined in Figure 4, obtaining values of 0.67 (I), 0.32 (II), 0.28 (III), and 0.07 (IV). This pronounced gradient demonstrates a clear dependence of FRB repeatability on the $\gamma$-$r$ parameter space. In particular, the stark contrast between the high-repetition Region I and the nearly non-repeating Region IV further quantifies the association between spectral morphology and repeat behavior. This result is consistent with the previous analysis based on density contours, reinforcing spectral morphology as a potential indicator for distinguishing repeaters.
Additionally, in Figure 4, we highlight two well-studied and highly active repeating FRB sources: FRB 20121102A and FRB 20201124A. Both have been monitored extensively with high sensitivity by the Arecibo telescope (Jahns et al., 2023) and FAST (Li et al., 2021; Xu et al., 2022b), yielding records of thousands of bursts, thereby providing critical insights into the origins of FRBs. Notably, both sources fall within the high-repetition-probability Region I defined in our framework, serving as compelling evidence for the robustness and predictive usefulness of this spectral partitioning. Except for a small number of atypical repeating bursts, this framework successfully assigns approximately 80% of known repeaters to the high-probability Region I. Overall, this probability map directly links FRB repeatability to spectral morphology parameters, offering a simple, physically interpretable prediction framework independent of prior labeling.
To assess the performance of the $\gamma$-$r$ framework in assigning repetition probability scores, we randomly withhold 10% of known repeaters (corresponding to 21 FRBs) and 10% of apparently non-repeating FRBs (49 FRBs) from the construction of the $\gamma$-$r$ framework, and build the framework using the remaining FRB sample. We then estimate the repetition probability scores for the held-out known repeaters.
Given the limited number of confirmed repeating FRBs in the current dataset, we repeat this procedure 100 times to evaluate the robustness of the framework under different random realizations. Figure 5 shows the distributions of the mean (left) and median (right) repetition probability scores for the held-out repeaters across 100 random subsampling realizations. While variations in the randomly selected subsets introduce modest fluctuations in the inferred scores, the mean repetition probability score averaged over all realizations is 0.65, generally consistent with the repetition probability density of Region I in the full $\gamma$-$r$ framework (Figure 4).
The median scores, which characterize the typical repetition probability of the held-out repeaters and are insensitive to a small number of extreme values, are strongly concentrated in the range 0.9-1.0 for the vast majority of realizations, with only a few cases extending into the intermediate-probability regime, consistent with the presence of a small number of repeaters with atypical spectral properties. This behavior indicates that, even when the framework is constructed from different data subsets, at least half of the held-out known repeaters are consistently assigned high repetition probabilities, demonstrating the robustness of the $\gamma$-$r$ framework.
Figure 5: The left and right panels show the distributions of the mean and median repetition probability scores, respectively, for the randomly held-out 10% known repeaters, obtained from 100 random subsampling realizations, with repetition probability scores estimated using the $\gamma$-$r$ framework.
For future FRBs newly detected by CHIME, their positions in this map can be readily determined using only $r$ and $\gamma$, enabling a preliminary estimation of their likelihood to exhibit repeating FRBs. This provides a practical and accessible tool for rapidly assessing repetition probability and prioritizing monitoring targets in future observations. Compared with machine learning classification methods that rely on multi-parameter dimensionality reduction and clustering, this physically based partitioning is faster, more intuitive, and yields robust, generalizable results.
We further combine the t-SNE results with the $\gamma$-$r$ parameter map to identify a set of FRB sources that have not yet been observed to repeat but exhibit statistical characteristics closely resembling those of known repeaters. Specifically, we first selected all non-repeating FRBs classified within the Repeater-like cluster and categorized them according to their quadrant positions in the $\gamma$-$r$ parameter space, with particular attention to those located in the high-repetition-probability Region I.
@@
IV Conclusion
In this work, we analyzed the repeatability characteristics of FRBs using an extended CHIME/FRB sample, integrating dimensionality reduction, clustering, feature-importance evaluation, and spectral-parameter-based partitioning. We further propose a simple and generalizable method that relies solely on spectral parameters to classify FRBs by their repeatability preference and to effectively estimate their repetition probabilities. The main conclusions are as follows: